OpenStudy (anonymous):

solve the system by substitution. 8x+y=17 x+4y=37

terenzreignz (terenzreignz):

All right, this is one of the more basic techniques in solving systems of linear equations, so pay attention ^_^ In the first equation, what I want you to do is try to get y alone on one side of the equals-sign... can you do that?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

It's simple... for example, in this (easier) case... x - y = 5 Suppose I want to get x alone on one side of the equation... well, no big deal, all I have to do is add +y to both sides to get.... x - y + y = 5 + y -y + y cancels out on the left, leaving me with... x = 5 + y And now x is alone on the left side. Now, can you get y alone on the left side in 8x + y = 17 ?